what is the answer for the equation f(x)=x^4-6x^3+11x^2+12x-26 divided by 3-2i
Answer by Edwin McCravy(20060)

f(x)=x^4-6x^3+11x^2+12x-26 divided by 3-2i


   3-2i |  1    -6    11    12    -26
        |____________________________

Bring down the 1

   3-2i |  1    -6      11    12    -26
        |                              
           1

Multiply the 1 by the 3-2i: 1(3-2i) = 3-2i.  Write that under the -6
like this:

   3-2i |  1    -6      11    12    -26
        |        3-2i                  
           1
  
Add -6 and 3-2i, getting -3-2i.  Put that underneath the line to
the right of the 1, like this:

   3-2i |  1    -6      11    12    -26
        |        3-2i                  
           1    -3-2i

Multiply the -3-2i by the 3-2i: (-3-2i)(3-2i) = -9+6i-6i+4i² 
= -9+4(-1) = -9-4 = -13.  Write than under the 11 like this:

   3-2i |  1    -6      11    12    -26
        |        3-2i  -13             
           1    -3-2i


Add 11 and -13, getting -2.  Put that underneath the line to
the right of the -3-2i, like this:

   3-2i |  1    -6      11    12    -26
        |        3-2i  -13             
           1    -3-2i   -2

Multiply the -2 by the 3-2i: -2(3-2i) =-6+4i.  Write that under the 12
like this:

   3-2i |  1    -6      11    12     -26
        |        3-2i  -13    -6+4i     
           1    -3-2i   -2     6+4i

Multiply the 6+4i by the 3-2i: (6+4i)(3-2i) = 18-12i+12i-8i² 
= 18-8(-1) = 18+8 = 26.  Write than under the -26 like this:

   3-2i |  1    -6      11    12     -26
        |        3-2i  -13    -6+4i   26
           1    -3-2i   -2     6+4i

Add and get 0 remainder

   3-2i |  1    -6      11    12     -26
        |        3-2i  -13    -6+4i   26
           1    -3-2i   -2     6+4i    0

We interpret that bottom line as coefficients of a
polynomial of one less degree than the original.  The
degree of the original is 4, so this will be of degree 3:

   x³ + (-3-2i)x² - 2x + (6+4i)

That's the answer. or you can factor out a negative
from the second coefficiens:

   x³ - (3+2i)x² - 2x + (6+4i)

Edwin